1. A patient needs the administration of 12 mCi F-FDG and the current concentration of F-FDG is 8.3 mCi/mL. How much volume should be prepared? A. 0.69 mL B. 0.95 mL C. 1.44 mL D. 2.03 mL 2. The weight of a patient is 134 lb and the physician has prescribed 0.14 mCi/kg for this procedure. How much activity should be prepared? A. 7.5 mOi B. 8.5 mCi C. 14.3 mCi D. 18.7 mCi 3. 1F-FLT is helpful in patients with cancer because the uptake mechanism demonstrates the rate of A. blood flow B. fatty acid metabolism C. tumor hypoxia D. cellular proliferation may be performed with all of the following PET T 4. Myocardial perfusion studie radiopharmaceuticals EXCEP A. 1F-FDG B. N-ammonia C. 1O-water D. Rb-chloride
Answer #1

Answer: ( C) 1.44 ml

The concentration available is 8.3 mci/kg.

The patient needs to administer 12mci.

Therefore the volume to be prepared will be= 12 ÷ 8.3= 1.44ml.

Answer:( B) 8.50

Solution:

1lb = 0.4535 kilograms

The patient weight is 134lb, so, 134× 0.4535 = 60.76 kg.

The drug prescribed 0.14mci/kg

Therefore, 60.76kg × 0.14mci= 8.50mci drug should be given.

Answer:( A) Blood flow

Answer: (C) O- water